The height of the students in a certain class is following normal distribution with mean height as 165 cm and standard deviation of 25 cm. There are 60 students in that class. Determine
i. The number of students whose height is more than 158 cm.
ii. The number of students whose height is lying between 155 and 172 cm.
1
Expert's answer
2020-06-15T19:21:29-0400

"i) Given \; that, μ=165, σ=25, n=60, then,\\
 P(x>158)=P(Z>\frac{158-165}{25})\\
=P(Z>-0.28)=0.5+P(0<Z<0.28)\\
=0.5+0.1103=0.6103\\
 number\;of\;students=60\times0.6103=36.63\approx37\\
ii) P(155<x<172)\\
=P(\frac{155-165}{25})<Z<\frac{172-165}{25})\\
=P(-0.4<Z<0.28)\\
=P(0<Z<0.4)+P(0<Z<0.28)\\
=0.1554+0.1103=0.2657\\
 number\;of\;students=60\times0.2657=15.94\approx16\\"